# smalltableaux

### v1

@@ I Introduction
In many of these applications, the goal is to extract information that does not depend on the order of the given qubits-that is, permutationally invariant (PI) information. This includes, for instance, quantum metrology and state tomography [1, 2, 3, 4, 5], where the goal is to use many copies of a quantum state to estimate an unknown parameter or reconstruct a classical description of the state. Another example discussed further below is qubit purification [6], where many noisy copies of an unknown pure state are combined to obtain one or more copies with higher fidelity. Other important examples of applications include density matrix exponentiation [7], quantum emulation [8], coherence distillation [9], and majority voting [10].
@@
We show that for a system of $n$ qubits, after approximately $2n\ln n$ SWAP tests, a sharp transition occurs: with high probability, no additional singlets are found, and the state of the remaining qubits converges to a fixed state within their totally symmetric subspace. More formally, we compare the output of this protocol with that of the so-called Schur transform [11, 12, 13], a powerful tool in quantum information with broad applications, which, in particular, has been commonly used for extracting PI information. We prove that for PI states, the trace distance between the output of the Schur transform and the protocol based on SWAP tests is bounded by $\epsilon>0$, provided that $T$, the number of SWAP tests, satisfies
@@
As an example of applications, we demonstrate the application of this protocol for state purification [6]. Namely, we apply it to $\rho^{\otimes n}$, i.e., $n$ copies of state $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$, where $|\psi\rangle$ is an unknown pure state and $0<p<1$, in order to obtain a single qubit with higher fidelity with $|\psi\rangle$. Cirac et al.[6] have shown how the optimal fidelity can be achieved using the Schur transform, and the fidelity of the output $\rho_{\text{opt}}(n)$ with the desired output state $|\psi\rangle$ is given by
@@
Our result then reveals that, up to an arbitrarily small additional error $\epsilon$, the same fidelity can be achieved using $2n\ln(n\epsilon^{-1})$ random SWAP tests, by simply returning one of the qubits that has not been discarded by the SWAP test. See Fig. 2 for an example, and Sec. VII for further details.
@@
We note that, prior to this work, Childs et al. [14] used sequential SWAP tests to implement qubit purification. However, in their approach, when a SWAP test identifies a pair of qubits in the triplet subspace, one qubit is retained and forwarded to the next level, and the other is discarded. This simplifies the analysis, as the remaining qubits are fully uncorrelated, but it leads to wasted resources and thus suboptimal fidelity as well as suboptimal gate complexity (See Sec.VII).
We also note that in the context of symmetry-protected quantum computation, Freedman et al. [15] showed in 2021 that the outcomes of random SWAP tests, referred to as triplet/singlet measurements in their work, determine the total angular momentum of a system of qubits. Building on this, Rudolph and Virmani [16] showed in 2023 that, starting from qubits in the maximally mixed state, one can, by repeatedly measuring angular momentum in this way, gradually grow a subset of qubits whose joint state is restricted to their totally symmetric subspace, namely, in the maximally mixed state on this subspace. (At each step, one performs random SWAP tests between the previously obtained subset and a new qubit in the maximally mixed state).
Our results reveal broader applications of random SWAP tests, namely for performing the Schur transform on PI states, which is useful, for example, in the context of state purification and tomography, and they provide more precise and tighter error analysis, uncovering a sharp threshold phenomenon around $T_{\ast}\approx 2n\ln n$.  [^1] Rudolph and Virmani [16] show that the convergence can be achieved to the maximally mixed state on the totally symmetric subspace in $\approx n^{2}$ steps. Freedman et al. [15] briefly discuss using random SWAP tests for measuring angular momentum $j$, as a useful subroutine in the so-called symmetry-protected quantum computation. They argue that “since spin is quantized, the convergence is actually faster than polynomial. Once the number of measurements is polynomially large compared to $S^{2}$, the convergence in accuracy becomes exponential.”
Figure 2: The fidelity of the output qubit of this purification protocol, with the desired ideal pure state, after $T$ SWAP tests. Initially the system contains $n=100$ copies of $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$. As the number of SWAP tests grows, the fidelity of the output state converges to the optimal fidelity given by Eq.(2). The vertical dotted line corresponds to $n\ln n$, which roughly indicates the order of the number of required SWAP tests. See Sec.VII for further details.
@@ II Setup
Consider a system of $n$ qubits with the total Hilbert space $(\mathbb{C}^{2})^{\otimes n}$. We are interested in the properties of quantum states that are independent of the order of the qubits. More formally, for any state $\sigma$, we are interested in the information that can be obtained from the PI state $\mathcal{P}(\sigma)$, where $\mathcal{P}$ is the quantum channel defined by applying a permutation on $n$ qubits, chosen uniformly at random from the group of all such permutations $\mathbb{S}_{n}$ (See Appendix A for further details). By applying such random permutations, any state $\sigma$ is projected to its symmetrized version $\mathcal{P}(\sigma)$. Hence, in the following, we often restrict our attention to the case of PI states.
@@ III Unitary Schur Sampling
For many applications in quantum information and computation, it is useful to physically realize the transformation that, for a given PI input state $\sigma$, measures $j$, i.e., returns the angular momentum value $j$ with probability $p_{j}$, along with the corresponding post-measurement state $\sigma_{j}$. In [17], this operation is called unitary Schur sampling.
In practice, for many applications, such as purification [6] and coherence distillation [9], rather than returning the post-measurement state $\sigma_{j}$ as a subsystem with dimension $2j+1$, it is often more convenient and useful to have an output system whose Hilbert space is of constant size. A natural and convenient way to achieve this is to encode the state $\sigma_{j}$ in the $(2j+1)$-dimensional totally symmetric subspace of $2j$ qubits. The remaining n2jn-2j qubits (which is always an even number) can be prepared in pure singlet states, i.e., in the state $|\xi\rangle=(|01\rangle-|10\rangle)/\sqrt{2}$ (As we see below, this choice is made to respect the $\mathrm{SU}(2)$ symmetry). That is, the goal is to realize the transformation
@@
where $\xi=|\xi\rangle\langle\xi|$, $\widetilde{\sigma}_{j}$ is an unnormalized density operator in the totally symmetric subspace of $2j$ qubits satisfying $\operatorname{Tr}(\widetilde{\sigma}_{j})=p_{j}$, and we have appended the system with an additional (classical) register that keeps track of the value of $j$ as $|j\rangle_{\text{reg}}$. [^2] Obviously, this register does not need to be a separate quantum device; it could be, for example, the controller of the experimental apparatus or simply the experimentalist’s memory! As we further explain below in Sec.VIII.1, if one is interested in obtaining state $\sigma_{j}$ encoded in $\lceil\log_{2}(2j+1)\rceil$ qubits, this can be achieved using an extra compression circuit, which converts $\widetilde{\sigma_{j}}/p_{j}$ to $\sigma_{j}$.
To explicitly write the matrix elements of $\widetilde{\sigma}_{j}$, we first consider a basis for the totally symmetric subspace of $2j$ qubits, namely, the states $|j,m\rangle$ for $m=-j,\dots,j$, the normalized eigenvectors of $\sum_{i=1}^{2j}\sigma_{z}^{(i)}/2$ in this subspace, also known as the Dicke states. In particular, $|j,m\rangle$ is the uniform superposition of all computational basis states in $\{|0\rangle,|1\rangle\}^{\otimes 2j}$ that contain $j+m$ qubits in state $|0\rangle$ and jmj-m qubits in state $|1\rangle$ (See Appendix A) for the explicit formula). Then,
@@
is a basis for a single copy of an irrep of $\mathrm{SU}(2)$ with angular momentum $j$ in $\mathbb{C}^{2j+1}\otimes\mathbb{C}^{m(n,j)}$. More precisely, with respect to this decomposition, these vectors span the subspace $\mathbb{C}^{2j+1}\otimes|\eta_{0}\rangle$, where $|\eta_{0}\rangle$ is a fixed state in $\mathbb{C}^{m(n,j)}$, the multiplicity subsystem associated to irrep $j$ (See Appendix A for further discussion).

### v2

# Optimal Qubit Purification and Unitary Schur Sampling via Random SWAP Tests

@@ I Introduction
In many of these applications, the goal is to extract information that does not depend on the order of the given qubits-that is, permutationally invariant (PI) information. This includes, for instance, quantum metrology and state tomography [helstrom1969quantum, holevo2011probabilistic, giovannetti2011advances, giovannetti2006quantum, TomographyHarrow], where the goal is to use many copies of a quantum state to estimate an unknown parameter or reconstruct a classical description of the state. Another example discussed further below is qubit purification [Cirac], where many noisy copies of an unknown pure state are combined to obtain one or more copies with higher fidelity. Other important examples of applications include density matrix exponentiation [densitylloyd], quantum emulation [marvian2016universal], coherence distillation [marvian2020coherence], and majority voting [majority].
@@
We show that for a system of $n$ qubits, after approximately $2n\ln n$ SWAP tests, a sharp transition occurs: with high probability, no additional singlets are found, and the state of the remaining qubits converges to a fixed state within their totally symmetric subspace. More formally, we compare the output of this protocol with that of the so-called Schur transform [weakchilds, SchurTransformHarrow, harrowthesis], a powerful tool in quantum information with broad applications, which, in particular, has been commonly used for extracting PI information. We prove that for PI states, the trace distance between the output of the Schur transform and the protocol based on SWAP tests is bounded by $\epsilon>0$, provided that $T$, the number of SWAP tests, satisfies
@@
As an example of applications, we demonstrate that this protocol for state purification [Cirac]. Namely, we apply it to $\rho^{\otimes n}$, i.e., $n$ copies of state $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$, where $|\psi\rangle$ is an unknown pure state and $0<p<1$, in order to obtain a single qubit with higher fidelity with respect to $|\psi\rangle$. Cirac et al. [Cirac] have shown how the optimal fidelity can be achieved using the Schur transform, and the fidelity of the output $\rho_{\text{opt}}(n)$ with the desired output state $|\psi\rangle$ is given by
@@
Our result then reveals that, up to an arbitrarily small additional error $\epsilon$, the same fidelity can be achieved using $2n\ln(n\epsilon^{-1})$ random SWAP tests, by simply returning one of the qubits that has not been discarded by the SWAP test. See Fig. 2 for an example, and Sec. LABEL:Sec:App:Pur for further details.
@@
We note that, prior to this work, Childs et al. [Childs2025streamingquantum] used sequential SWAP tests to implement qubit purification. However, in their approach, when a SWAP test identifies a pair of qubits in the triplet subspace, one qubit is retained and forwarded to the next level, and the other is discarded. This simplifies the analysis, as the remaining qubits are fully uncorrelated, but it leads to wasted resources and thus suboptimal fidelity as well as suboptimal gate complexity (see Sec.LABEL:Sec:App:Pur).
We also note that in the context of symmetry-protected quantum computation, Freedman et al. [freedman2021symmetry] showed in 2021 that the outcomes of random SWAP tests, referred to as triplet/singlet measurements in their work, determine the total angular momentum of a system of qubits. Building on this, Rudolph and Virmani [rudolph] showed in 2023 that, starting from qubits in the maximally mixed state, one can, by repeatedly measuring angular momentum in this way, gradually grow a subset of qubits whose joint state is restricted to their totally symmetric subspace, namely, in the maximally mixed state on this subspace. (At each step, one performs random SWAP tests between the previously obtained subset and a new qubit in the maximally mixed state).
Our results reveal broader applications of random SWAP tests, namely for performing the Schur transform on PI states, which is useful, for example, in the context of state purification and tomography, and they provide more precise and tighter error analysis, uncovering a sharp threshold phenomenon around $T_{\ast}\approx 2n\ln n$.  [^1] Rudolph and Virmani [rudolph] show that the convergence can be achieved to the maximally mixed state on the totally symmetric subspace in $\approx n^{2}$ steps. Freedman et al. [freedman2021symmetry] briefly discuss using random SWAP tests for measuring angular momentum $j$, as a useful subroutine in the so-called symmetry-protected quantum computation. They argue that “since spin is quantized, the convergence is actually faster than polynomial. Once the number of measurements is polynomially large compared to $n^{2}$, the convergence in accuracy becomes exponential.” We also note that, in the special case of $n=3$ qudits, it was previously observed in [li2024optimal] that unitary Schur sampling can be achieved via sequential SWAP tests.
Figure 2: The fidelity of the output qubit of this purification protocol, with the desired ideal pure state, after $T$ SWAP tests. Initially the system contains $n=100$ copies of $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$. As the number of SWAP tests grows, the fidelity of the output state converges to the optimal fidelity given by Eq.(2). The vertical dotted line corresponds to $n\ln n$, which roughly indicates the order of the number of required SWAP tests. See Sec.LABEL:Sec:App:Pur for further details.
@@ II Setup
Consider a system of $n$ qubits with the total Hilbert space $(\mathbb{C}^{2})^{\otimes n}$. We are interested in the properties of quantum states that are independent of the order of the qubits. More formally, for any state $\sigma$, we are interested in the information that can be obtained from the PI state $\mathcal{P}(\sigma)$, where $\mathcal{P}$ is the quantum channel defined by applying a permutation on $n$ qubits, chosen uniformly at random from the group of all such permutations $\mathbb{S}_{n}$ (See Appendix LABEL:App:SU(2) for further details). By applying such random permutations, any state $\sigma$ is projected to its symmetrized version $\mathcal{P}(\sigma)$. Hence, in the following, we often restrict our attention to the case of PI states.
@@ III Unitary Schur Sampling
For many applications in quantum information and computation, it is useful to physically realize the transformation that, for a given PI input state $\sigma$, measures $j$, i.e., returns the angular momentum value $j$ with probability $p_{j}$, along with the corresponding post-measurement state $\sigma_{j}$. In [cervero2024memory], this operation is called unitary Schur sampling.
In practice, for many applications, such as purification [Cirac] and coherence distillation [marvian2020coherence], rather than returning the post-measurement state $\sigma_{j}$ as a subsystem with dimension $2j+1$, it is often more convenient and useful to have an output system whose Hilbert space is of constant size. A natural and convenient way to achieve this is to encode the state $\sigma_{j}$ in the $(2j+1)$-dimensional totally symmetric subspace of $2j$ qubits. The remaining n2jn-2j qubits (which is always an even number) can be prepared in pure singlet states, i.e., in the state $|\xi\rangle=(|01\rangle-|10\rangle)/\sqrt{2}$ (As we see below, this choice is made to respect the $\mathrm{SU}(2)$ symmetry). That is, the goal is to realize the transformation
@@
where $\xi=|\xi\rangle\langle\xi|$, $\widetilde{\sigma}_{j}$ is an unnormalized density operator in the totally symmetric subspace of $2j$ qubits satisfying $\operatorname{Tr}(\widetilde{\sigma}_{j})=p_{j}$, and we have appended the system with an additional (classical) register that keeps track of the value of $j$ as $|j\rangle_{\text{reg}}$. [^2] Obviously, this register does not need to be a separate quantum device; it could be, for example, the controller of the experimental apparatus or simply the experimentalist’s memory! As we further explain below in Sec.LABEL:Sec:compress, if one is interested in obtaining state $\sigma_{j}$ encoded in $\lceil\log_{2}(2j+1)\rceil$ qubits, this can be achieved using an extra compression circuit, which converts $\widetilde{\sigma_{j}}/p_{j}$ to $\sigma_{j}$.
To explicitly write the matrix elements of $\widetilde{\sigma}_{j}$, we first consider a basis for the totally symmetric subspace of $2j$ qubits, namely, the states $|j,m\rangle$ for $m=-j,\dots,j$, the normalized eigenvectors of $\sum_{i=1}^{2j}\sigma_{z}^{(i)}/2$ in this subspace, also known as the Dicke states. In particular, $|j,m\rangle$ is the uniform superposition of all computational basis states in $\{|0\rangle,|1\rangle\}^{\otimes 2j}$ that contain $j+m$ qubits in state $|0\rangle$ and jmj-m qubits in state $|1\rangle$ (see Appendix LABEL:App:SU(2) for the explicit formula). Then,
@@
is a basis for a single copy of an irrep of $\mathrm{SU}(2)$ with angular momentum $j$ in $\mathbb{C}^{2j+1}\otimes\mathbb{C}^{m(n,j)}$. More precisely, with respect to this decomposition, these vectors span the subspace $\mathbb{C}^{2j+1}\otimes|\eta_{0}\rangle$, where $|\eta_{0}\rangle$ is a fixed state in $\mathbb{C}^{m(n,j)}$, the multiplicity subsystem associated to irrep $j$ (See Appendix LABEL:App:SU(2) for further discussion).